A chemist, has one 30 litre bottle of 15% hydrochloric acid and one 30 litre bottle of 90% hydrochloric acid. He mixes 20 litres of 60% hydrochloric acid and then pours 5 L of that solution into the bottle containing the 90% hydrochloric acid. How strong is the acid in that bottle now?
Answer by richwmiller(17219)

20*.6+.9*5=x
.6*20=12
.9*5=4.5
x=16.5
20+5=25
We know it should be between 60 and 90 but closer to 60
16.5/25=.66 =66%